40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 9 01:36:28 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OLtrend.png
    Reply  Mon May 12 22:42:19 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OL-FB.pngarms_140512.pdf
       Reply  Mon May 12 22:52:59 2014, Jenne, Summary, SUS, Optical Lever Filters are all different OpLevFilters.pdf
       Reply  Tue May 13 13:27:58 2014, Steve, Summary, SUS, Optical Lever Servos setting table OLsums.png
          Reply  Tue May 13 17:03:05 2014, Steve, Summary, SUS, ETMX oplev laser replaced 
             Reply  Tue May 13 17:31:32 2014, Jenne, Summary, SUS, ETMX oplev laser replaced: New oplev gains set OLerr_13May2014.pdfOLfb_13May2014.pdf
                Reply  Tue May 13 22:55:57 2014, Jenne, Summary, SUS, ETMX oplev: cleanup 
             Reply  Wed May 14 10:04:06 2014, Steve, Summary, SUS, oplev laser and temperature TempRules.png
                Reply  Tue May 20 09:12:25 2014, Steve, Summary, SUS, ETMX oplev qpd board ETMXOLQPD121.jpgETMXOLQPDface121.jpgETMXolSUM.png
                   Reply  Tue May 20 15:54:39 2014, rana, Summary, SUS, ETMX oplev qpd board 
Message ID: 9942     Entry time: Mon May 12 22:42:19 2014     In reply to: 9934     Reply to this: 9943   9946
Author: rana 
Type: Summary 
Category: SUS 
Subject: Optical Lever QPD Sum trends: they're almost all too weak 

For some reason or another, I decided that we should see if the optical lever servos were injecting too much noise into the test masses. The ITMs are much worse than the ETMs and I am afeared that they might be making the main noise for our arms in the 20-40 Hz region. Jenne is checking up on these feedback loops to see what's up.

To estimate the actuator gains of the mirrors, I turned on 1 count drives from LSC/CAL oscillators into the LSC drives of each test mass at the frequencies shown in the plot with the resulting peaks showing up in in POX/Y with the single arm locks in red. I will leave these going permanently, but with 0.1 count ampltiudes; we need to make it so in the scripts.

Attachment 1: OL-FB.png  106 kB  | Hide | Hide all
OL-FB.png
Attachment 2: arms_140512.pdf  9 kB  Uploaded Mon May 12 23:57:30 2014  | Hide | Hide all
arms_140512.pdf
ELOG V3.1.3-